Write out the equation -23 times the number -78 is equal to -71 less than the number
Answer by VictoryStar(6)

-23 times a number is -23x.
-78 is fine on it's own.
Equal to -71 less than the number would be x(the number) - (less than ) -71.
Combine and receive -23x - 78 = x--71
Since two minuses make a plus, it becomes -23x - 78 = x + 71.